# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 4.1 Martingale problem
$+\int_{0}^{t-s}e^{-(q_{i}+\beta_{i})r}\mathbb{E}_{\chi_{i}(s),i}\bigg[\beta_{i}g_{i}(\mathbf{u}(t-s-r,\chi_{i}(r),i))+\sum_{j\neq i}q_{ij}\mathbf{u}(t-s-r,U_{ij}+\chi_{i}(r),j)\bigg]\mathrm{d}r.$ (4.5)
@@
$\qquad+\int_{s}^{t}\!\!\!e^{-(q_{i}+\beta_{i})(w-s)}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\mathbf{u}(t-w,\chi_{i}(w),i))+\sum_{j\neq i}q_{ij}\mathbf{u}(t-w,U_{ij}+\chi_{i}(w),j)\bigg]\mathrm{d}w.$ (4.8)
@@
$\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(x+\widehat{X}_{u}(t),J_{u}(t))\bigg]$
$=\mathbb{E}_{x,i}\left[\Phi_{\theta}(\widehat{\chi}_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\Phi_{\theta}(\widehat{\chi}_{i}(t-s),i))+\sum_{j\neq i}q_{ij}\Phi_{\theta}(U_{ij}+\widehat{\chi}_{i}(t-s),j)-(q_{i}+\beta_{i})\Phi_{\theta}(\widehat{\chi}_{i}(t-s),i)\bigg]\mathrm{d}s.$
Due to (4.2), the latter is equal to $\Phi_{\theta}(x,i)$. Therefore, $\mathbf{E}_{x,i}\Big[\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(x+\widehat{X}_{u}(t),J_{u}(t))\Big]=\Phi(x,i)$ is a constant for every $t$. It follows from the Markov property that $\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(\widehat{X}_{u}(t),J_{u}(t))$ is a martingale under $\mathbf{P}_{x,i}$. ∎
@@ 4.2 Proof of existence of travelling waves
Suppose $\theta\in(0,\bar{\theta})$ with $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. Define a function $\Phi_{\theta}\colon\mathbb{R}\times\mathcal{I}\to[0,1]$ by $(x,i)\mapsto\Phi_{\theta}(x,i)=\mathbf{E}_{x,i}\big[e^{-W_{\theta}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}W_{\theta}(\infty)}\right]$. Then $\Phi_{\theta}$ is a travelling wave solution with speed $\frac{\lambda(\theta)}{\theta}$.
@@
This yields that $\mathbf{E}_{x,i}\big[e^{-W_{\theta}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}W_{\theta}(\infty)}\right]$ and $\Phi_{\theta}(x,i)$ is well defined. Moreover, it follows from Theorem 1.4 that when $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$, $W_{\theta}(\infty)$ is non-degenerate and $\mathbf{P}_{0,i}(W_{\theta}(\infty)=0)=\mathbf{q}_{i}$. Thus we have
@@
$u(t,x,i):=\Phi_{\theta}\Big(x-\frac{\lambda(\theta)}{\theta}t,i\Big)=\mathbf{E}_{0,i}\left[\exp\left\{-e^{\lambda(\theta)t-\theta x}W_{\theta}(\infty)\right\}\right]=\mathbf{E}_{x,i}\left[\exp\left\{-e^{\lambda(\theta)t}W_{\theta}(\infty)\right\}\right].$ (4.13)
@@
$=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{s}}\mathbf{E}_{X_{u}(s),J_{u}(s)}\left[\exp\left\{-e^{\lambda(\theta)(t-s)}W_{\theta}^{(u)}(\infty)\right\}\right]\bigg]=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{s}}u(t-s,X_{u}(s),J_{u}(s))\bigg].$ (4.15)
@@
Suppose $\sum_{k\geq 1}k(\log k)^{2}\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. Then the function $\Phi_{\theta^{*}}(x,i)=\mathbf{E}_{x,i}\big[e^{-Z_{\theta^{*}}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}Z_{\theta^{*}}(\infty)}\right]$ is a travelling wave solution with speed $\lambda^{\prime}(\theta^{*})=\frac{\lambda(\theta^{*})}{\theta^{*}}$.
@@
Using the argument similar to the proof of Lemma 4.3, we have that $\Phi_{\theta^{*}}$ is well defined and $\Phi_{\theta^{*}}\in\mathcal{T}_{1}$. Recall that $Z_{\theta^{*}}(t)=e^{-\lambda(\theta^{*})t}\sum_{u\in\mathcal{N}_{t}}e^{-\theta^{*}X_{u}(t)}\left[V_{J_{u}(t)}(\theta^{*})(X_{u}(t)+\lambda^{\prime}(\theta^{*})t)-V^{\prime}_{J_{u}(t)}(\theta^{*})\right]$ and $Z_{\theta^{*}}(\infty)=\lim_{t\rightarrow\infty}Z_{\theta^{*}}(t)$ almost surely under $\mathbf{P}_{x,i}$. Therefore, we know that under $\mathbf{P}_{x,i}$,
@@
By Lemmas 4.3 and 4.4, it remains to show the non-existence of travelling waves with speed $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}$. The proof is an extension of classical arguments (see e.g. [Har-FKPP]). By Corollary 1.5 the velocity of the leftmost particle is $\lambda^{\prime}(\theta^{*})$. As $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}=\lambda^{\prime}(\theta^{*})$, we have $\mathbf{P}_{x,i}$-a.s.
@@
$\Phi(x,i)$ $=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{t}}\Phi(X_{u}(t)+\rho t,J_{u}(t))\bigg]=\mathbf{E}_{x,i}\bigg[\left(\mathbbm{1}_{\left\{\mathscr{S}^{c}\right\}}+\mathbbm{1}_{\left\{\mathscr{S}\right\}}\right)\prod_{u\in\mathcal{N}_{t}}\Phi(X_{u}(t)+\rho t,J_{u}(t))\bigg]$ (4.20)
@@ 4.3 Proof of uniqueness of travelling waves
We now prove the uniqueness of travelling waves, for which we further assume that the branching MAP is spectrally negative; this assumption means that, for all $i\in\mathcal{I}$, the Lévy process $\chi_{i}$ has no positive jumps (its Lévy measure satisfies $\Lambda_{i}(0,\infty)=0$) and $\mathbb{P}(U_{ij}\leq 0)=1$ for all $i,j\in\mathcal{I}$. In this proof, we follow the general ideas of [KYPRIANOU04, YangRen11].
@@
When a particle crosses this barrier, it is stopped immediately. According to [Sato99, Theorem 46.2], a Lévy process without positive jumps that crosses an upper barrier from below must do so continuously. Consequently, a particle governed by such a process is stopped upon hitting the barrier. Let $C(x,\rho_{\theta})$ denotes the random collection of particles stopped at the barrier, then is a stopping line, as it satisfies the fundamental property that, if $u\in C(x,\rho_{\theta})$, then $v\notin C(x,\rho_{\theta})$ for all $v\prec u$; see [Chauvin91, Jagers89]. Then we have the following properties.
@@
  • For $x<y$ and any $u\in C(y,\rho_{\theta})$, there exists a unique $v\in C(x,\rho_{\theta})$ such that $v\prec u$. This follows from the first point and the fact that when the particles hit the barrier. Let $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}$ be the natural filtration generated by ancestral types and spatial paths receding from particles at the moment they hit $\Gamma^{(x,\rho_{\theta})}$ (see [Chauvin91, Jagers89] for precise definition). Therefore, for $x<y$, $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\subset\mathcal{F}_{\Gamma^{(y,\rho_{\theta})}}$.
@@
    We use $\sigma_{u}^{x}$ to denote the stopping time when the particle $u$ hit the barrier $\Gamma^{(x,\rho_{\theta})}$. By the strong Markov branching property on stopping lines (see [Chauvin91, Neveu87] for branching Brownian motions and [Kyprianou99] for branching Lévy processes), we know that given $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}$, the processes $(\#C_{j}(y,\rho_{\theta}):j\in\mathcal{I},y>x;\mathbf{P}_{(X_{u}(\sigma_{u}^{x}),\sigma_{u}^{x}),J_{u}(\sigma_{u}^{x})})$ for $u\in\mathcal{N}_{t}$ are independent. Therefore, $(\#C_{j}(x,\rho_{\theta}):j\in\mathcal{I};\mathbf{P}_{0,i})$ satisfies the definition on [Athreya1968, Page 200] and forms a continuous-time branching process indexed by $x\geq 0$.
@@
Furthermore, by the Kesten-Stigum theorem (see, for example, [Georgii03, Theorem 2.1]), we know that
@@
We next define
$\widetilde{W}_{x}(\rho_{\theta}):=\sum_{u\in C(x,\rho_{\theta})}e^{-\theta X_{u}(\sigma_{u}^{x})-\lambda(\theta)\sigma_{u}^{x}}V_{J_{u}(\sigma_{u}^{x})}(\theta)=\sum_{i\in\mathcal{I}}\#C_{i}(x,\rho_{\theta})V_{i}(\theta)e^{-\theta x},\qquad x\geq 0.$ (4.26)
@@
$\mathbf{E}_{0,i}$ $\left[M_{\infty}(z,\rho_{\theta})|\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\right]=\lim_{t\rightarrow\infty}\mathbf{E}_{0,i}\left[M_{t}(z,\rho_{\theta})|\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\right]$ (4.38)
$=\lim_{t\rightarrow\infty}\bigg(\prod_{u\in C(x,\rho_{\theta}),\sigma_{u}^{x}\leq t}\Phi_{\theta}(z+X_{u}(\sigma_{u}^{x})+\rho_{\theta}\sigma_{u}^{x},J_{u}(\sigma_{u}^{x}))\times\prod_{u\in\mathcal{N}_{t},\sigma_{u}^{x}>t}\Phi_{\theta}(z+X_{u}(t)+\rho_{\theta}t,J_{u}(t))\bigg)$ (4.39)
@@
Similarly to the arguments for supercritical speed regime, we know that $(\#C_{i}(x,\rho_{\theta^{*}}):i\in\mathcal{I})_{x\geq 0}$ forms a supercritical continuous-time branching process where $x$ plays the role of time. Again, by the Kesten-Stigum theorem ([Georgii03, Theorem 2.1]), there is a non-negative vector $\pi^{\theta^{*}}=(\pi^{\theta^{*}}_{i}:i\in\mathcal{I})$ with $\langle\pi^{\theta^{*}},1\rangle=1$, such that for all $i\in\mathcal{I}$, $\mathbf{P}_{0,i}(\lim_{x\rightarrow\infty}\#C_{j}(x,\rho_{\theta^{*}})/\#C(x,\rho_{\theta^{*}})=\pi^{\theta^{*}}_{j}\mid\mathscr{S})=1$. Therefore,

### v2

@@ 4.1 Martingale problem
$+\int_{0}^{t-s}e^{-(q_{i}+\beta_{i})r}\mathbb{E}_{\chi_{i}(s),i}\bigg[\beta_{i}g_{i}(\mathbf{u}(t-s-r,\chi_{i}(r),i))+\sum_{j\neq i}q_{ij}\mathbf{u}(t-s-r,U_{ij}+\chi_{i}(r),j)\bigg]\mathrm{d}r.$
@@
$\qquad+\int_{s}^{t}\!\!\!e^{-(q_{i}+\beta_{i})(w-s)}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\mathbf{u}(t-w,\chi_{i}(w),i))+\sum_{j\neq i}q_{ij}\mathbf{u}(t-w,U_{ij}+\chi_{i}(w),j)\bigg]\mathrm{d}w.$
@@
$\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(x+\widehat{X}_{u}(t),J_{u}(t))\bigg]=\mathbb{E}_{x,i}\left[\Phi_{\theta}(\widehat{\chi}_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\Phi_{\theta}(\widehat{\chi}_{i}(t-s),i))$
$+\sum_{j\neq i}q_{ij}\Phi_{\theta}(U_{ij}+\widehat{\chi}_{i}(t-s),j)-(q_{i}+\beta_{i})\Phi_{\theta}(\widehat{\chi}_{i}(t-s),i)\bigg]\mathrm{d}s.$
Due to (4.2), the latter is equal to $\Phi_{\theta}(x,i)$. Therefore,
$\mathbf{E}_{x,i}\Big[\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(x+\widehat{X}_{u}(t),J_{u}(t))\Big]=\Phi(x,i)$ (4.8)
is a constant for every $t$. It follows from the Markov property that $\prod_{u\in\mathcal{N}_{t}}\Phi_{\theta}(\widehat{X}_{u}(t),J_{u}(t))$ is a martingale under $\mathbf{P}_{x,i}$. ∎
@@ 4.2 Proof of existence of travelling waves
Suppose $\theta\in(0,\bar{\theta})$ with $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\mathbbm{1}_{\left\{\beta_{j}m_{j}>0\right\}}\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. Define a function $\Phi_{\theta}\colon\mathbb{R}\times\mathcal{I}\to[0,1]$ by $(x,i)\mapsto\Phi_{\theta}(x,i)=\mathbf{E}_{x,i}\big[e^{-W_{\theta}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}W_{\theta}(\infty)}\right]$. Then $\Phi_{\theta}$ is a travelling wave solution with speed $\frac{\lambda(\theta)}{\theta}$.
@@
This yields that $\mathbf{E}_{x,i}\big[e^{-W_{\theta}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}W_{\theta}(\infty)}\right]$ and $\Phi_{\theta}(x,i)$ is well defined. Moreover, when $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$, it follows from Theorem 1.4 that $W_{\theta}(\infty)$ is non-degenerate and $\mathbf{P}_{0,i}(W_{\theta}(\infty)=0)=\mathbf{q}_{i}$. Thus we have
@@
$u(t,x,i):$ $=\Phi_{\theta}\Big(x-\frac{\lambda(\theta)}{\theta}t,i\Big)=\mathbf{E}_{0,i}\left[\exp\left\{-e^{\lambda(\theta)t-\theta x}W_{\theta}(\infty)\right\}\right]$ (4.12)
$=\mathbf{E}_{x,i}\left[\exp\left\{-e^{\lambda(\theta)t}W_{\theta}(\infty)\right\}\right].$ (4.13)
@@
$=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{s}}\mathbf{E}_{X_{u}(s),J_{u}(s)}\left[\exp\left\{-e^{\lambda(\theta)(t-s)}W_{\theta}^{(u)}(\infty)\right\}\right]\bigg]$ (4.15)
$=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{s}}u(t-s,X_{u}(s),J_{u}(s))\bigg].$ (4.16)
@@
Suppose $\mathbbm{1}_{\left\{\beta_{j}m_{j}>0\right\}}\sum_{k\geq 1}k(\log k)^{2}\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. Then the function $\Phi_{\theta^{*}}(x,i)=\mathbf{E}_{x,i}\big[e^{-Z_{\theta^{*}}(\infty)}\big]=\mathbf{E}_{0,i}\left[e^{-e^{-\theta x}Z_{\theta^{*}}(\infty)}\right]$ is a travelling wave solution with speed $\lambda^{\prime}(\theta^{*})=\frac{\lambda(\theta^{*})}{\theta^{*}}$.
@@
Using the argument similar to the proof of Lemma 4.3, we have that $\Phi_{\theta^{*}}$ is well defined and $\Phi_{\theta^{*}}\in\mathcal{T}_{1}$. Recall that
$Z_{\theta^{*}}(t)=e^{-\lambda(\theta^{*})t}\sum_{u\in\mathcal{N}_{t}}e^{-\theta^{*}X_{u}(t)}\left[V_{J_{u}(t)}(\theta^{*})(X_{u}(t)+\lambda^{\prime}(\theta^{*})t)-V^{\prime}_{J_{u}(t)}(\theta^{*})\right]$ (4.18)
and $Z_{\theta^{*}}(\infty)=\lim_{t\rightarrow\infty}Z_{\theta^{*}}(t)$ almost surely under $\mathbf{P}_{x,i}$. Therefore, we know that under $\mathbf{P}_{x,i}$,
@@
By Lemmas 4.3 and 4.4, it remains to show the non-existence of travelling waves with speed $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}$. The proof is an extension of classical arguments (see e.g. [24]). By Corollary 1.5 the velocity of the leftmost particle is $\lambda^{\prime}(\theta^{*})$. As $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}=\lambda^{\prime}(\theta^{*})$, we have $\mathbf{P}_{x,i}$-a.s.
@@
$\Phi(x,i)$ $=\mathbf{E}_{x,i}\bigg[\prod_{u\in\mathcal{N}_{t}}\Phi(X_{u}(t)+\rho t,J_{u}(t))\bigg]$ (4.22)
$=\mathbf{E}_{x,i}\bigg[\left(\mathbbm{1}_{\left\{\mathscr{S}^{c}\right\}}+\mathbbm{1}_{\left\{\mathscr{S}\right\}}\right)\prod_{u\in\mathcal{N}_{t}}\Phi(X_{u}(t)+\rho t,J_{u}(t))\bigg]$ (4.23)
@@ 4.3 Proof of uniqueness of travelling waves
We now prove the uniqueness of travelling waves, for which we further assume that the branching MAP is spectrally negative; this assumption means that, for all $i\in\mathcal{I}$, the Lévy process $\chi_{i}$ has no positive jumps (its Lévy measure satisfies $\Lambda_{i}(0,\infty)=0$) and $\mathbb{P}(U_{ij}\leq 0)=1$ for all $i,j\in\mathcal{I}$. In this proof, we follow the general ideas of [37, 56].
@@
When a particle crosses this barrier, it is stopped immediately. According to [52, Theorem 46.2], a Lévy process without positive jumps that crosses an upper barrier from below must do so continuously. Consequently, a particle governed by such a process is stopped upon hitting the barrier. Let $C(x,\rho_{\theta})$ denotes the random collection of particles stopped at the barrier, then is a stopping line, as it satisfies the fundamental property that, if $u\in C(x,\rho_{\theta})$, then $v\notin C(x,\rho_{\theta})$ for all $v\prec u$; see [15, 31]. Then we have the following properties.
@@
  • For $x<y$ and any $u\in C(y,\rho_{\theta})$, there exists a unique $v\in C(x,\rho_{\theta})$ such that $v\preceq u$. This follows from the first point and the fact that when the particles hit the barrier. Let $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}$ be the natural filtration generated by ancestral types and spatial paths receding from particles at the moment they hit $\Gamma^{(x,\rho_{\theta})}$ (see [15, 31] for precise definition). Therefore, for $x<y$, $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\subset\mathcal{F}_{\Gamma^{(y,\rho_{\theta})}}$.
@@
    We use $\sigma_{u}^{x}$ to denote the stopping time when the particle $u$ hit the barrier $\Gamma^{(x,\rho_{\theta})}$. By the strong Markov branching property on stopping lines (see [15, 46] for branching Brownian motions and [36] for branching Lévy processes), we know that given $\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}$, the processes $(\#C_{j}(y,\rho_{\theta}):j\in\mathcal{I},y>x;\mathbf{P}_{(X_{u}(\sigma_{u}^{x}),\sigma_{u}^{x}),J_{u}(\sigma_{u}^{x})})$ for $u\in\mathcal{N}_{t}$ are independent. Therefore, $(\#C_{j}(x,\rho_{\theta}):j\in\mathcal{I};\mathbf{P}_{0,i})$ satisfies the definition on [5, Page 200] and forms a continuous-time branching process indexed by $x\geq 0$.
@@
Furthermore, by the Kesten-Stigum theorem (see, for example, [21, Theorem 2.1]), we know that
@@
We next define for $x\geq 0$
$\widetilde{W}_{x}(\rho_{\theta}):=\sum_{u\in C(x,\rho_{\theta})}e^{-\theta X_{u}(\sigma_{u}^{x})-\lambda(\theta)\sigma_{u}^{x}}V_{J_{u}(\sigma_{u}^{x})}(\theta)=\sum_{i\in\mathcal{I}}\#C_{i}(x,\rho_{\theta})V_{i}(\theta)e^{-\theta x}.$ (4.29)
@@
$\mathbf{E}_{0,i}\left[M_{\infty}(z,\rho_{\theta})|\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\right]$ (4.41)
$=\lim_{t\rightarrow\infty}\mathbf{E}_{0,i}\left[M_{t}(z,\rho_{\theta})|\mathcal{F}_{\Gamma^{(x,\rho_{\theta})}}\right]$ (4.42)
$=\lim_{t\rightarrow\infty}\bigg(\prod_{u\in C(x,\rho_{\theta}),\sigma_{u}^{x}\leq t}\!\!\!\Phi_{\theta}(z\!+\!X_{u}(\sigma_{u}^{x})\!+\!\rho_{\theta}\sigma_{u}^{x},J_{u}(\sigma_{u}^{x}))\prod_{u\in\mathcal{N}_{t},\sigma_{u}^{x}>t}\Phi_{\theta}(z\!+\!X_{u}(t)\!+\!\rho_{\theta}t,J_{u}(t))\bigg)$ (4.43)
@@
Similarly to the arguments for supercritical speed regime, we know that $(\#C_{i}(x,\rho_{\theta^{*}}):i\in\mathcal{I})_{x\geq 0}$ forms a supercritical continuous-time branching process where $x$ plays the role of time. Again, by the Kesten-Stigum theorem ([21, Theorem 2.1]), there is a non-negative vector $\pi^{\theta^{*}}=(\pi^{\theta^{*}}_{i}:i\in\mathcal{I})$ with $\langle\pi^{\theta^{*}},1\rangle=1$, such that for all $i\in\mathcal{I}$, $\mathbf{P}_{0,i}(\lim_{x\rightarrow\infty}\#C_{j}(x,\rho_{\theta^{*}})/\#C(x,\rho_{\theta^{*}})=\pi^{\theta^{*}}_{j}\mid\mathscr{S})=1$. Therefore,